OpenStudy (khantahmina):

Hi! Please help with the question below. Will be fanned don't worry and medals will be given :) Help with number 19 below

OpenStudy (tashywuzhere):

put ur question

OpenStudy (khantahmina):

Right here being attached

OpenStudy (joannablackwelder):

That first piece looks good. :-) Do you know the form of a circle function?

OpenStudy (khantahmina):

yeah the equation of a circle yeah (x-h)^2 +(y-k)^2=r^2

OpenStudy (khantahmina):

@JoannaBlackwelder

OpenStudy (joannablackwelder):

Great! And the piece that is circular has what center and radius?

OpenStudy (khantahmina):

r u referring to the firrst pg

OpenStudy (joannablackwelder):

Yeah

OpenStudy (joannablackwelder):

That is what the question is asking for, right?

OpenStudy (khantahmina):

well i did all that . i need help with number 19 the last question which is the 3rd pg

OpenStudy (joannablackwelder):

Right, but f(x) is referring back to the first page, is it not? Or is there another f(x) given?

OpenStudy (khantahmina):

no ur right. i think its from the first pg bc then otherwise theres no f(x) only g(x) n etc

OpenStudy (joannablackwelder):

Cool. :-)

OpenStudy (khantahmina):

soo i have one equation x-5 5<x<6

OpenStudy (joannablackwelder):

Right, and can you write the circle equation with the center and radius that we have given in the pic?

OpenStudy (khantahmina):

(x-h)^2 +(y-k)^2 =r^2 (x-0)^2 + (y-0)^2=r^2 (h,k)=(0,0) (x,y)= any random point (0,3) (0-0)^2 +(3-0)^2=r^2 0+9=r^2 9=r^2 radius is 3 equation of this circle is (x-0)^2 +(y-0)^2= 9

OpenStudy (khantahmina):

but this is a semi circle

OpenStudy (joannablackwelder):

Great! Right. To write that in a function form, we will need to simplify and solve for y.

OpenStudy (joannablackwelder):

Can you do that step?

OpenStudy (khantahmina):

uggggg :)

OpenStudy (joannablackwelder):

You're doing great! :-)

OpenStudy (khantahmina):

wait so i have a question

OpenStudy (joannablackwelder):

Sure

OpenStudy (khantahmina):

bc this is a semi circle right, should it be (x-0)^2 +(y-0)^2= 9

OpenStudy (joannablackwelder):

We'll get to making it just the semi circle in just a sec, once we have it solved for y. For now, it is the whole circle. Good question. :-)

OpenStudy (khantahmina):

oh thanks im working on to making it to y=...

OpenStudy (khantahmina):

ooh i got it. y=3-x

OpenStudy (joannablackwelder):

Close, but not quite. Let me show you.

OpenStudy (joannablackwelder):



OpenStudy (joannablackwelder):